For a 6% coupon bond with a 1,200 face value selling at par with 2 years to maturity calculate the rate of return for this year if you sell the bond 1 year later and if its yield to maturity at the beginning of this year is 6% and at the end of this year the interest rate unexpectedly rises to 10%.
Answer #1

Given about a bond,

Coupon rate = 6%

Facr value = $1200

So, coupon = 6% of 1200 = $72

Years to maturity = 2 years

YTM = 6%

Since bond is selling at Par, so price of bond now is = face value

So price now = $1200

after 1 year, interest rate rises to 10%

So, now the bond price = 1272/1.1 = $1156.364

So bond is purchased at $1200

sold at $1156.364 and a coupon of $72 is received at end of year 1

rate of return = (Final price + coupon - initial price)/initial price = (1156.364 + 72 - 1200)/1200 = 2.364%